OpenStudy (anonymous):

Natalie works in a toy shop and earns $43 per day. She earns an extra $3 for each toy she sells. If Natalie wants to earn at least $70 per day, which inequality shows the minimum number of toys, n, that she should sell? 43 + 3n ≥ 70, so n ≥ 9 43 + 3n ≤ 70, so n ≤ 9 43 + 3n ≥ 70, so n ≥ 24 43 + 3n ≤ 70, so n ≤ 24

OpenStudy (anonymous):

@VeritasVosLiberabit last one :P

OpenStudy (anonymous):

the inequality will need to be greater than or equal to 70 because that is the minimum required to earn 70$ per day

OpenStudy (anonymous):

ok so i can already remove the less then ones

OpenStudy (anonymous):

Correct. Now you need to analyze the difference between the two remaining A and C

OpenStudy (anonymous):

mmhh

OpenStudy (anonymous):

The difference being the n values associated with the inequalities. This requires performing a bit of algebra by solving for n. Treat the inequality sign just like an equals sign.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would I get to the answer then?

OpenStudy (anonymous):

im like the worst with these types of questions sorry:/

OpenStudy (anonymous):

Solve the equation using algebra just like a normal algebraic equation.

OpenStudy (anonymous):

i mean how would I do that im the worst with these questions...

OpenStudy (anonymous):

\[43+3n \ge 70\] solve for n

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

man i dont know :(

OpenStudy (anonymous):

@VeritasVosLiberabit i have no idea :/

OpenStudy (anonymous):

here @KennyVanGemert solve this then: 43+3n=70

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

Is it just the sign that is throwing you off?

OpenStudy (anonymous):

idk i think its trying to figure the N out

OpenStudy (anonymous):

and the sign

OpenStudy (anonymous):

23?

OpenStudy (anonymous):

tell me where you are having difficulty. You have to isolate n on one side of the equation and then isolate the numbers on the otherside.

OpenStudy (anonymous):

can you solve this? x+4=5?

OpenStudy (anonymous):

the n

OpenStudy (anonymous):

1+4=5

OpenStudy (anonymous):

ah you are solving by inspection. can you solve this one then? 2x+6= 34?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ok show me

OpenStudy (anonymous):

i think its 28 because 28+6=34

OpenStudy (anonymous):

Ok so you are still soving by inspection. You should be solving these algebraically

OpenStudy (anonymous):

im the worst at these idk really how to do that

OpenStudy (anonymous):

I will show you 

OpenStudy (anonymous):

hmm let me look at that a few more times hang on

OpenStudy (anonymous):

the key is to try an isolate the value of x by performing algebraic operations on both sides of the equation until you get a value.

OpenStudy (anonymous):

oh ok i think i get it let me try it with the question now

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

43+3n=70 just posting this so i can see it

OpenStudy (anonymous):

43+3n-3=67

OpenStudy (anonymous):

then u have the 43 and 67

OpenStudy (anonymous):

You want to try and isolate n. So what is the first step you should start with?

OpenStudy (anonymous):

um subtracting the 3 from the 70

OpenStudy (anonymous):

the three is attached to 3n since n is mutiplied by 3 so that isn't where to start

OpenStudy (anonymous):

ohhh ok so u start with the 43

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ohhh ok let me try this now

OpenStudy (anonymous):

so then you do 70-43 and get 27?

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

now you should have 3n=27. What next?

OpenStudy (anonymous):

umm lemme look

OpenStudy (anonymous):

u would do hm

OpenStudy (anonymous):

hm?

OpenStudy (anonymous):

im trying to think xD

OpenStudy (anonymous):

would u do division? i dont think u do but would u?

OpenStudy (anonymous):

9!!!!!

OpenStudy (anonymous):

That's actually right. Why don't you think it's correct/

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

i was dividing wrong

OpenStudy (anonymous):

soo it would be A

OpenStudy (anonymous):

there you go. So then the answer is clearly A

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

wow thankk u soooooooooooooooooooooo much if i had 100$ i would give it to u thank u soooooooooooooooooooooooooooooooooooooooooooooooo much

OpenStudy (anonymous):

ok im gonna go do the easy ones thank u again

OpenStudy (anonymous):

you're welcome.